OpenStudy (anonymous):

Write a matrix equation to solve for the variables. Write your solution as an ordered triple (x,y,z). x+y-z=14 4x-y+5z=-22 2x+2y-3z=35

OpenStudy (anonymous):

LOLOLLOLOLLOLOLLOLOL

OpenStudy (anonymous):

\[ \left[\begin{matrix} 1&1&-1\\ 4&-1&5\\ 2&2&-3 \end{matrix}\right]\cdot \left[\begin{matrix} x\\y\\z \end{matrix}\right]= \left[\begin{matrix} 14\\-22\\35 \end{matrix}\right]\\ \text{We can create an augmented matrix:}\\ \left[\begin{array}{ccc|c} 1&1&-1&14\\ 4&-1&5&-22\\ 2&2&-3&35 \end{array}\right]\\ \text{Then use elimination:}\\ \left[\begin{array}{ccc|c} 1&1&-1&14\\ 0&-5&9&-78\\ 0&0&-1&7 \end{array}\right]\\ \left[\begin{array}{ccc|c} 1&1&0&7\\ 0&-5&0&-15\\ 0&0&1&-7 \end{array}\right]\\ \left[\begin{array}{ccc|c} 1&0&0&4\\ 0&1&0&3\\ 0&0&1&-7 \end{array}\right]\\ \text{And we have our result: }(4,3,-7) \]

OpenStudy (anonymous):

I WATCHED TE MATRIX just find neo